A 900g stone is pushed along a tarmac by a horizontal force of 20N. A frictional force of 8N opposes the motion. What is the acceleration given to the stone?
1
Expert's answer
2020-09-10T14:39:59-0400

"F=ma\to a=\frac{F}{m}=\frac{20-8}{0.9}=13.3(m/s^2)"